OpenStudy (anonymous):

Which expression is equivalent to cos 2 pi for all values of pi?

OpenStudy (anonymous):

well you should just check out some trigonmetric formulas for cos, i will give you a hint, look for one with cos^2 x

OpenStudy (anonymous):

\[\cos 2 \pi=1\]

OpenStudy (anonymous):

Wouldn't it be cos^2pi – 1 ?

OpenStudy (anonymous):

\[2\cos^2\pi-1 \implies 1-\sin^2\pi \implies \cos^2\pi-\sin^2\pi\]

OpenStudy (anonymous):

yes with the 2 before

OpenStudy (anonymous):

2cos^2pi-1

OpenStudy (anonymous):

yes. cos(2pie) is equivalent to cos(0) which is equal to "1"

OpenStudy (anonymous):

cos2pi + sin2pi cos2pi – 1 1 – 2 sin2 pi 2 sinpi cospi are the options

OpenStudy (anonymous):

sin(2pi)=0 which makes first option as "cos2pi +0= cos2pi= 1". Therefore first option is the right one.

OpenStudy (anonymous):

1,3 are correct

OpenStudy (anonymous):

yup. 1 and 3.

OpenStudy (anonymous):

Oh, i see ! thank you :)

OpenStudy (anonymous):

i actually have one more question..

OpenStudy (anonymous):

post in a new open question tab

OpenStudy (anonymous):

i did :)